# Representational Analysis of Binding in Large Language Models

### v1

@@ 4.4 Results: Activation Steering on BI Subspace
Inspired by the research on Activation Steering (AS) Turner et al. (2023), we apply an AS method to verify the importance of the BI subspace on LM’s binding behaviour. Specifically, we use the following Equation 2 to extract a subspace steering vector $\mathbf{s}_{0\rightarrow bi}$, which is proposed to swap BI from 00 to $bi$, where $n$ is the number of target entities, $\mathbf{x}_{bi,l}^{i}$ represents the activation of entity $e_{i}$ from layer $l$, and its BI is $bi$. We intervene via Equation 3 and assume that by adding $\mathbf{s}_{0\rightarrow bi}$ to the original activation $\mathbf{x}_{0,l}$, we can increase the LD and the proportion of the attribute $a_{bi}$. Figure 5 shows the results on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). (Appendix (§A.7) shows the results on other datasets) These results indicate that AS can achieve the similar tendency as the direct value intervention mentioned in Section (§4.3). For instance, adding $\mathbf{s}_{0\rightarrow 3}$, which is used to swap BI form 00 to $3$, can increase the LD of $a_{3}$ and its proportion in the predicted answers. The consistent tendency with the results of the direct subspace editing, shown Figure 3 and Figure 4, further illustrates that the discovered subspace contains the BI information, and more importantly, it plays an important role when the model performs in-context binding computation.
@@ 4.5 Binding Subspace and Position
Table 3: Simplified expression of original inputs and the one modified with pseudo relation, which is proposed to equalize PI for PCA analysis, where $a_{0}^{p1}\,r\,e_{0}^{p2}$ represents a relation such as “the apple is in Box C”, and $e_{0}^{p2}$ denotes an entity with BI of 00 and PI of $p2$, $e_{2}^{p5}\ r^{-1}\ ?$ denotes the query on entity $e_{1}$, such as “Box C contains the”.
Input (original)
$a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{0}^{p1}\ r^{-1}\ ?$
Input (with interjection)
@@
In this section, we discuss the relationship between the BI subspace and Positional Information (PI), namely the $postion\_ids$ of input tokens. As mentioned in Section (§4.3), the discovered subspace stores the BI information, therefore direct intervention on it can swap the answer of LM. However, one counter hypothesis is that the subspace is not used for storing BI information but the PI of attributes, thus the direct intervention merely change the PI so that swap the answer. Regarding the relationship between BI and PI, Feng and Steinhardt (2023) found that even when PI of attributes is swapped, the model still makes correct predictions, thus confirming the independence between BI and PI. Based on this finding, we go one step further and illustrate the independence between the BI subspace and PI. To prove the independence, we create three datasets, the first one is by extending the original dataset with pseudo relation, as shown in Table 3, the second one is by inserting a sequence of interjection(e.g., “ah, ah, …”), as listed in Table 4, and the third one is by prefixing the original dataset with filler words(e.g., “It can be seen that”).
@@
New Dataset with Pseudo Relation. In Table 3, $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ refers to a pseudo relation, which is a fixed expression, such as “the PC is in Box Z”, applied to adjust the PI while keeping the BI. For instance, in Table 3, adding one or two $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ before $a_{1}\,r\,e_{1}$ (i.e., $C^{\prime}_{2}$ and $C^{\prime}_{1}$) does not affect the BI of $e_{1}$ but its PI, because $e_{1}$ is still the second unique entity from the left (i.e., BI$=1$), but its PI is $p3$ and $p5$ respectively. Using the pseudo relation, we create the data in a manner that the target entity (e.g., $e_{1}^{p5}$ and $e_{2}^{p5}$) to extract activation have the same PI, such as $C^{\prime}_{1}$ and $C^{\prime}_{2}$ in Table 3.
@@
We apply the method mentioned in Section (§2) on the set of activations $\{\overrightarrow{e_{1}^{p5}},\overrightarrow{e_{2}^{p5}},...\}$, where $\overrightarrow{e_{1}^{p5}}$ denotes the activation of $e_{1}$ in $e_{1}^{p5}\ r^{-1}\ ?$, so as to capture the BI difference and exclude the PI difference, because they share the same PI (i.e., $P5$) but different BI (i.e., $1$, $2$, …). Then we compare its BI subspace with the original one (e.g., $\{\overrightarrow{e_{1}^{p2}},\overrightarrow{e_{2}^{p5}},...\}$) to analyze how the distribution of BI subspace changes after removing the PI variance. Figure 6 visualizes the BI subspace distribution, where the light colored points denote the original distribution, and the dark ones are from the new one with equalized PI. We can observe that after removing the PI difference, the distribution is still similar to the original one that there is a clearly visible direction along which BI increases. This illustrates that our PCA based method can capture BI information, that is, along the direction of BI-PC, and it does not causally depend on PI. See Appendix (§A.5) for our further analysis on how the context of binding affects the distribution.
New Dataset with Interjections. The dataset is created by inserting a sequence of interjections after the first attribute entity pair, as illustrated in Table 4. Since there is no BI information in the interjection (e.g., $j^{p3}$), adding it only changes the PI of its following entities and attributes. We set the number of interjections as that its last PI is larger than the last PI of its original input (e.g., $pi>p5$ in Table 4).
Based on this dataset, we conduct the same intervention on its BI subspace, as mentioned in Section (§4.3). The counter argument is that the subspace only captures PI, and the intervening step only changes the PI information. Specifically, adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$ can change its PI from $p1$ to $p3$, which is the PI of $e_{1}$, and its attribute is $a_{1}$, as shown in Table 4, thereby the LM swaps the answer from $a_{0}$ to $a_{1}$. If it is true, then the same intervention will not change the answer on the new dataset, because following the counter argument, after adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$, its PI becomes $p3$, and $p3$ is the PI of $j^{p3}$. The LM thus would not select $a_{1}$ as its answer. However, the results on Figure 7 shows that the subspace intervention on the new dataset achieves similar results as the original one, as shown in Figure 3 and Figure 4, proving the counter argument wrong and indicating the independence between BI subspace and PI.
@@
New Dataset with Filler Words. The dataset is created by adding Filler Words(FW) with various length, such as “OK”, “I see that” and “There is no particular reason”, in front of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), as shown in Table 5. Since the length (i.e., the number of tokens) of FW directly affects the PI of its following entities and attributes without affecting their BI, we take the length as the measure of intervention on PI and apply Spearman’s rank correlation $\rho$ to calculate the correlation between the length (denoted as PI) and the discovered BI-PC. Figure 8 shows $\rho$ between PI and BI-PC as well as between BI and BI-PC. We can observe that BI-PC has high $\rho$ with BI but almost zero $\rho$ with PI, indicating that the discovered BI-PC is dependent on BI but independent on PI. Therefore, the BI-PC does not simply encode absolute token position.
@@
Figure 7: Logit Difference and Logit Flip for activation patching on the dataset of “r: is_in” with interjections. Appendix (§A.10) shows the results on other dataset with the same interjection based modification.
@@
4.6 Binding Subspace and Relational Information
@@
“I see apple, cup and hair, somewhere else there is Box E, the bell, bus, toy, cookie and Box F are scattered around, ”
@@
In order to study the relationship between BI-PC and Relational Information(RI), we create an alternative dataset by converting relational expression of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024) into non-related one. Specifically, we prepare a set of non-related expression templates and randomly select one to replace the original expression of relation as shown in Table 6. We can observe that the template could make a target EA pair (e.g., “apple” and “Box E”) semantically unrelated, but retain their Ordering Information(OI) (i.e., “apple”->“bell”, and “Box E”->“Box F”).
The BI mechanism Feng and Steinhardt (2023) notes that a LLM represents binding through the consistency of BI. According to the mechanism, we adopt the correlation score between BI-PC of attributes (e.g., BI-PC of “apple”) and of their corresponding entities (e.g., BI-PC of“Box E”) to represent their binding strength. We select Spearman’s rank correlation $\rho$ as the correlation score and present the $\rho$ of the non-related dataset as well as the related one in the Figure 9.
We can observe that there is strong correlation between BI-PC of attributes and of their corresponding entities in the original data, indicating that LLM applies the correlation of BI-PC to encode binding. We can also observe that $\rho$ of non-related dataset is lower than the related (i.e., original) one, indicating that the BI-PC contains a certain degree of RI so that removing RI can decrease the $\rho$. However, even there is no RI but OI in the non-related dataset, the BI-PC of attributes still has high correlation with their co-occurring entities, indicating that BI-PC not only stores RI, but also encodes OI, and the latter seems to be the main and fundamental factor that is utilized by the LLM to determine binding, because intervention on BI-PC can causally change the behaviour of LLM (as discussed in Section (§4). Therefore, we hypothesize that the BI-PC primarily encodes the OI of entities and attributes that is used as a prototype of BI to causally determine binding.
@@
4.7 Results on Fine-Tuned LM
Kramár et al. (2024); Kim et al. (2024) claim that the code fine-tuned LM, such as Float-7B Prakash et al. (2024) outperforms the pretrained LM on the entity tracking task Kim and Schuster (2023); Prakash et al. (2024). Since the code fine-tuned LM performs well on the entity tracking task that requires the BI subspace based computation, we hypothesize that BI subspace also exists in the code fine-tuned LM and the intervention along BI-PC will causally affect the model output. To prove the hypothesis, we conduct the intervention on Float-7B and show results in Figure 10. We found that the BI subspace based intervention on Float-7B achieves the similar results as on Llama2-7B, indicating that the BI subspace not only exists in the pretrained LM but also in the fine-tuned one. In addition, adding the same step value (i.e., $v$) on Float-7B will achieve higher LD value than Llama2-7B, indicating that the code fine-tuned LM is more sensitive to the BI subspace based intervention. For instance, the maximum LD of $a_{4}$ in the former is around $10$, and it is $2$ times larger than the one in the latter, which is around $5$. This might partially explains why the code fine-tuned LM performs better than the original one, because code fine-tuning might enhance the function of BI subspace so that it is more sensitive on the intervention and more effective on the in-context entity tracking task.
Figure 10: Logit Difference and Logit Flip for activation patching on the entity tracking dataset (i.e., r: is_in) on Float-7B. See Appendix (§A.8) for the results on other datasets.
@@ 5 Related Work
Linear Representation Recent research found that sequence models trained only on next token prediction linearly represent various semantic concepts including Othello board positions Li et al. (2022); Nanda et al. (2023), the truth value of assertions Marks and Tegmark (2023), sentiment Tigges et al. (2023), and numeric values such as elevation, population, birth year, and death year Gurnee and Tegmark (2023); Heinzerling and Inui (2024). Continuing this line of research, in this work, we discover that LMs such as Llama-2 can also linearly encode BI, because there is a linear direction that primarily encodes BI in the activations. See Appendix (§A.1) for other related work.
@@ 6 Conclusion and Future Work
In this work, we study the in-context binding, a fundamental skill underlying many complex reasoning and natural language understanding tasks. We provide a novel view of the Binding ID mechanism introduced by Feng and Steinhardt (2023) that there exists a low-rank subspace in the hidden state (or activation) of LMs that primarily encodes the ordering information and which is used as the prototype of BIs to causally determine binding. Our future work includes: 1. the analysis of BI subspace in a more realistic setting; 2. the study of interaction between in-context binding and factual knowledge learned from pretraining; 3. BI subspace based mechanistic analysis.
Limitation
The limitations of our research include the following points. Firstly, we only analyze BI subspace on the attribute prediction task, but not on the entity inference task (i.e., given an attribute to infer its entity). Secondly, we lack the analysis on how predicate (or relation) affect the BI subspace, and how the results of BI-subspace based intervention differ with the type of predicate. Thirdly, although we use a publicly available entity tracking dataset, it is still a synthesized dataset. Therefore, for uncovering how LMs bind and track entity in reality, it is necessary to analyze the BI subspace on a real world dataset. The last but not lest, we only analyze binding from the perspective of representation and localize BI subspace. However, we have not answered what is the mechanism that generates the subspace and what is the circuit that utilizes the subspace for binding.
@@
The existing dataset Kim and Schuster (2023); Prakash et al. (2024) and LMs (i.e., Llama2-7B, Float-7B and Llama3-8B) are applied according to their intended research purpose. The synthetic datasets we adopted in this work are automatically created by strictly following the rule (or pattern) of the existing dataset, where the entities and attributes are sampled from a pool of wide variety of one-token names and concepts. Therefore, there is no ethical concern on human annotation bias and semantic biases. The datasets and code will be publicly available to ensure the reproducibility of our experiments.

### v3

@@ 4.4 Results: Activation Steering on OI Subspace
Inspired by the research on Activation Steering (AS) Turner et al. (2023), we apply an AS method to verify the importance of the OI subspace on LM’s binding behaviour. Specifically, we use the following Equation 2 to extract a subspace steering vector $\mathbf{s}_{0\rightarrow bi}$, which is proposed to swap BI from 00 to $bi$, where $n$ is the number of target entities, $\mathbf{x}_{bi,l}^{i}$ represents the activation of entity $e_{i}$ from layer $l$, and its BI is $bi$. We intervene via Equation 3 and assume that by adding $\mathbf{s}_{0\rightarrow bi}$ on the original activation $\mathbf{x}_{0,l}$, we can increase the LD and the proportion of the attribute $a_{bi}$. Figure 5 shows the results on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). (Appendix (§A.8) shows the results on other datasets) These results indicate that AS can achieve the similar tendency as the direct value intervention mentioned in Section (§4.3). For instance, adding $\mathbf{s}_{0\rightarrow 3}$, which is supposed to swap BI from 00 to $3$, can increase the LD of $a_{3}$ and its proportion in the predicted answers. The consistent tendency with the results of the direct subspace editing, shown Figure 3 and Figure 4, further illustrates that the discovered OI subspace can causally affect BI information and binding behaviour. Therefore, OI subspace plays an important role when LMs perform in-context binding computation.
@@ 4.5 OI Subspace and Position
Table 3: Simplified expression of original inputs and the one modified with pseudo relation, which is proposed to equalize PI for PCA analysis, where $a_{0}^{p1}\,r\,e_{0}^{p2}$ represents a relation such as “the apple is in Box C”, and $e_{0}^{p2}$ denotes an entity with OI of 00 and PI of $p2$, $e_{2}^{p5}\ r^{-1}\ ?$ denotes the query on entity $e_{1}$, such as “Box C contains the”.
@@
In this section, we discuss the relationship between the OI subspace and Positional Information (PI), which is namely the $postion\_ids$ of input tokens. As mentioned in Section (§4.3), the discovered subspace can causally determine BI information. Therefore, direct intervention on the subspace can swap the answer of a LM. However, one counter hypothesis is that the subspace is not used for storing OI information but the PI of attributes, and thus the direct intervention merely changes the PI of answer token so that the LM applies the new PI to generate corresponding token. In other words, the OI space (or OI-PC) might have high correlation with PI but not OI.
To prove the independence between OI subspace and PI, we create three datasets, the first one is by extending the original dataset with pseudo relation, as shown in Table 3, the second one is by prefixing the original dataset with filler words(e.g., “It can be seen that”) and third one is by inserting a sequence of interjection(e.g., “ah, ah, …”), as listed in Table 12.
@@
New Dataset with Pseudo Relation. In Table 3, $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ refers to a pseudo relation, which is a fixed expression, such as “the PC is in Box Z”. The pseudo relation is applied to adjust the PI while keeping the OI. For instance, in Table 3, adding one or two $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ before $a_{1}\,r\,e_{1}$ (i.e., $C^{\prime}_{2}$ and $C^{\prime}_{1}$) does not affect the OI of $e_{1}$ but its PI, because $e_{1}$ is still the second unique entity from the left (i.e., its ordering index 0I$=1$), but its PI is $p3$ and $p5$ respectively. Using the pseudo relation, we create the data in a manner that the target entity for activation analysis (e.g., $e_{1}^{p5}$ and $e_{2}^{p5}$) have the same PI but different OI, such as $C^{\prime}_{1}$ and $C^{\prime}_{2}$ in Table 3.
@@
We apply the method mentioned in Section (§2) on the set of activations $\{\overrightarrow{e_{1}^{p5}},\overrightarrow{e_{2}^{p5}},...\}$, where $\overrightarrow{e_{1}^{p5}}$ denotes the activation of $e_{1}$ in query (i.e., $e_{1}^{p5}\ r^{-1}\ ?$), so as to capture the OI difference and exclude the PI difference, because they share the same PI (i.e., $P5$) but different OI (i.e., $1$, $2$, …). Then we compare its OI subspace with the original one (e.g., $\{\overrightarrow{e_{1}^{p2}},\overrightarrow{e_{2}^{p5}},...\}$) to analyze how the distribution of OI subspace changes after removing the PI variance. Figure 6 visualizes the OI subspace distribution, where the light colored points denote the original distribution, and the dark ones are from the new one with equalized PI. We can observe that after removing the PI difference, the distribution is still similar to the original one that there is a clearly visible direction along which OI increases. This illustrates that our PCA based method can capture OI information, that is, along the direction of OI-PC, and it does not causally depend on PI.
@@
New Dataset with Filler Words. The dataset is created by adding Filler Words (FW) with various length, such as “OK”, “I see that” and “There is no particular reason”, in front of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), as shown in Table 4. Since the length (i.e., the number of tokens) of FW directly changes the PI of its following entities and attributes without affecting their OIs, we take the length as the measure of intervention on PI and apply Spearman’s rank correlation $\rho$ to calculate the correlation between the length (denoted as PI) and the OI-PC value. Figure 7 shows $\rho$ between PI and OI-PC as well as between OI and OI-PC. We can observe that OI-PC has high $\rho$ with OI but almost zero $\rho$ with PI, indicating that the discovered OI-PC is highly correlated with OI information but independent on PI. Therefore, the OI-PC does not simply encode absolute token position. See Appendix (§A.11) for the analysis on the New Dataset with Interjections.
@@
4.6 Consistency of OI Subspace
The BI mechanism Feng and Steinhardt (2023) mentions that a LM represents a related EA pair through the consistency of their BI information. This naturally raises a research question that if there is the consistency between OI-PC of entities (e.g., OI-PC of “Box Z” in Sample 1) and of their corresponding attributes  [^4] The OI-PC of attribute is extracted from the query of attribute such as “(context) The coffee is in”. (e.g., OI-PC of“coffee”).
To analyze the consistency of OI-PC between a related EA pair, we prepare four alternative datasets with various binding pattern as shown in Table 6. We use the distance on OI-PC to measure consistency and show partial result of mutual OI-PC based distance in Figure 8. We can observe that potentially related EA pairs (e.g., “E0_0” and “A0_0”) tend to have lower OI-PC based distance than arbitrary pairs in all binding patterns, indicating that to some extent, OI-PC based distance could be seen as an important feature to represent binding. To further illustrate its significance, we attempt to classify related EA pairs only relying on their OI-PC based distances. Specifically, we search an optimal threshold value from a development set, and which is applied to classify potentially related EA pairs in a testing set. The results are shown in Figure 9. We can observe that the performance of the OI-PC (i.e., “PC1”) based distance is significantly better than other PCs across all binding patterns, indicating that comparing to other PCs, the OI-PC could be used to compute binding information.
4.7 OI Subspace and Relatedness
@@
“I see apple, somewhere else there is Box E, the bell and Box F are scattered around, …”
@@
Input (7A-7E)
“A0 is in E0, A1 is in E1, A2 is in E2, A3 is in E3, A4 is in E4, A5 is in E5, A6 is in E6.”
Input (7A-3E)
“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E1, A5 is in E2, A6 is in E2.”
Input (7A-2E)
“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E1, A5 is in E1, A6 is in E1.”
Input (7A-5E)
“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E2, A5 is in E3, A6 is in E4.”
Table 6: Simplified expression of the datasets with various binding pattern, where “7A-3E” represents the pattern containing 7 attributes and 3 entities.
(a) 7A-7E
(b) 7A-3E
(c) 7A-2E
(d) 7A-5E
@@
Binding of a EA pair means that the EA pair is bound by a binding relation such as property and location Treisman (1996). In turn, there is no binding when the EA pair is unrelated in its context. This raises a research question that if the correlation between OI-PC of attributes (e.g., OI-PC of “apple”) and of their corresponding entities (e.g., OI-PC of “Box E”) represents the relatedness namely the existence of a relation.
In order to uncover the relationship between OI-PC and the relatedness, we create an alternative dataset by converting relational expression of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024) into non-related one. Specifically, we prepare a set of non-related expression templates and randomly select one to replace the original expression of relation as shown in Table 5. We can observe that the template could make a target EA pair (e.g., “Box E” and “apple”) semantically unrelated but retain their OI (e.g., the OI of “apple” and “bell” are still 00 and $1$ respectively). We select Spearman’s rank correlation $\rho$ as the correlation metric and compare the $\rho$ of the non-related dataset with the related one in the Figure 10.
We can observe that $\rho$ of non-related dataset is slightly lower than the related (i.e., original) one, indicating that the OI-PC might contain limited relational information so that removing it can marginally decrease the $\rho$. However, there is still strong correlation between the non-related (or non-bound) entity attribute pair, indicating that the OI-PC primarily encodes the OI information but not binding information specifically the information of binding relation.
[…]
@@ 5 Related Work
Linear Representation Recent research found that sequence models trained only on next token prediction linearly represent various semantic concepts including Othello board positions Li et al. (2022); Nanda et al. (2023), the truth value of assertions Marks and Tegmark (2023), sentiment Tigges et al. (2023), and numeric values such as elevation, population, birth year, and death year Gurnee and Tegmark (2023); Heinzerling and Inui (2024). Continuing this line of research, in this work, we discover that multiple LMs such as Llama2 can also linearly encode OI along a OI increasing direction in the activations.
See Appendix (§A.1) for additional related work.
@@ 6 Conclusion and Future Work
In this work, we study the in-context binding, a fundamental skill underlying many complex reasoning and natural language understanding tasks. We provide a novel view of the Binding ID mechanism introduced by Feng and Steinhardt (2023) that there exists a low-rank subspace in the hidden state (or activation) of LMs that primarily encodes the ordering information and which is used as the prototype of BIs to causally determine binding. Our future work includes: 1. the analysis of OI subspace in a more realistic setting; 2. the study of interaction between in-context binding and factual knowledge learned from pretraining; 3. OI subspace based mechanistic analysis.
Acknowledgement
This work was supported by JST CREST Grant Number JPMJCR20D2 and JSPS KAKENHI Grant Number 21K17814. We are grateful to the anonymous reviewers for their constructive comments.
Limitations
The limitations of our research include the following points: 1. We only analyze OI subspace on the attribute prediction task, but not on the entity inference task (i.e., given an attribute to infer its entity); 2. Although PCA based analysis is empirically proven to be effective, we lack the theoretical analysis on why PCA could capture OI subspace; 3. We lack the analysis on how predicate (or relation) affect the OI subspace, and how the results of OI subspace based intervention differ with the type of predicate; 4. Although we use a publicly available entity tracking dataset, it is still a synthesized dataset. Therefore, for uncovering how LMs bind and track entity in reality, it is necessary to analyze the binding via a real world dataset; 5. We only analyze the activations in the query part instead of in the context part, thus this work can not explain how LMs encode OI in the context and how it is used for OI encoding in query part and how it contributes the binding computation; 6. We infer the change of BI information via the result of model output but the interaction between OI and BI information is not directly observed. This research thus lacks the mechanistic interpretation on the interaction between them; 7. We only analyze binding from the perspective of representation and localize OI subspace. However, we have not answered what is the mechanism that generates the subspace and what is the circuit that utilizes the subspace for binding.
@@ Acknowledgement
The existing dataset Kim and Schuster (2023); Prakash et al. (2024) and LMs (i.e., Llama2-7B, Float-7B, Llama3-8B, Qwen1.5-7B and Pythia-6.9B) are applied according to their intended research purpose. The synthetic datasets we adopted in this work are automatically created by strictly following the rule (or pattern) of the existing dataset, where the entities and attributes are sampled from a pool of wide variety of one-token names and concepts. Therefore, there is no ethical concern on human annotation bias and semantic biases.